# Securing Federated Learning against Backdoor Threats with Foundation Model Integration

### v2

@@ I Introduction
However, incorporating FMs into FL systems (FM-FL) introduces new complexities and exacerbates existing threats. Recent studies [15, 16] have identified a novel backdoor mechanism in FL arising from the interaction between FL and FM. Consider a classic FM-FL scenario where the FM on the server generates synthetic data that mirrors the client data distribution for use in global model aggregation. An attacker can exploit FM vulnerabilities to embed a backdoor into the synthetic data. This backdoor is then planted into the global model during the aggregation stage using the poisoned synthetic dataset. Subsequently, the backdoor is transmitted to all FL clients as the poisoned global model is distributed, thereby compromising the integrity of the entire system.
@@
It is critical to develop an effective defense strategy that addresses both novel and classic backdoor threats within the FL framework. A common attack pattern in both types of attacks is the abnormally large activations in the hidden layers of the compromised model when processing manipulated inputs [30, 19]. To mitigate this, a feasible approach is to impose upper bounds on internal activations. These activation bounds aim to constrain activations to reasonable limits, effectively mitigating malicious behavior on poisoned samples while preserving the model’s functionality on normal samples. The optimization of these bounds leverages synthetic data, maintaining local data independence in FL. This process is performed after global model fusion on the server, with the original model parameters fixed, ensuring both functionality and defense efficiency. Hence, this method provides an effective defense against backdoor attacks while remaining well-suited to the FL framework.
@@ II-A Backdoor Attacks in FL
Classic vs. Novel Attack Mechanisms: Classic backdoor threats stem from a subset of compromised clients and propagate to the central server, with malicious updates identified as anomalies among client updates. In contrast, novel attacks originate on the server, embedding threats into the global model and affecting all clients uniformly, making anomalies undetectable.

### v3

@@ I Introduction
However, incorporating FMs into FL systems (FM-FL) introduces new complexities and exacerbates existing threats. Recent studies [16, 17, 18, 15] have identified a novel backdoor mechanism in FL arising from the interaction between FL and FM. Consider a classic FM-FL scenario where the FM on the server generates synthetic data that mirrors the client data distribution for use in global model aggregation. An attacker can exploit FM vulnerabilities to embed a backdoor into the synthetic data. This backdoor is then planted into the global model during the aggregation stage using the poisoned synthetic dataset. Subsequently, the backdoor is transmitted to all FL clients as the poisoned global model is distributed, thereby compromising the integrity of the entire system.
@@
It is critical to develop an effective defense strategy that addresses both novel and classic backdoor threats within the FL framework. A common attack pattern in both cases is the abnormally large activations in hidden layers of compromised models when processing manipulated inputs [32, 21]. To mitigate this, a feasible approach is to impose upper bounds on internal activations. These bounds constrain activations to reasonable limits, effectively mitigating malicious behavior on poisoned samples while preserving model functionality on normal ones. The optimization of these bounds leverages synthetic data, maintaining local data independence in FL. This process is applied after global model fusion on the server, with original model parameters fixed, ensuring both functionality and defense efficiency. Hence, this method provides effective defense against backdoor attacks while remaining well-suited to the FL framework.
@@ II-A Backdoor Attacks in FL
Classic vs. Novel Attack Mechanisms: Classic backdoor threats stem from compromised clients and propagate to the server, with malicious updates identified as anomalies among client updates. In contrast, novel attacks originate on the server, embedding threats into the global model and affecting all clients uniformly, making anomalies undetectable.